OpenStudy (anonymous):

Find dy/dx by implicit differentiation. sqxy=x^7y+61

OpenStudy (richyw):

\[\sqrt{xy}=x^7y+61\]?

OpenStudy (anonymous):

yes

OpenStudy (shamil98):

Show your attempt at finding dy/dx. Remember sqrt(xy) = (xy)^1/2

OpenStudy (anonymous):

1/2(xy)^-1/2 (xy'+y) 7x^6y+x^7y'

OpenStudy (anonymous):

i got it thank you